how do I solve these problems:
y = x - 600 =

y = 1.25X + 350 =

Answer by stanbon(75887)

how do I solve these problems:
y = x - 600 =
y = 1.25X + 350 =
----------------------
substitute to get:
x-600 = 1.25x+350
0.25x = -950
x = -3000
=====================
Substitute to solve for "y":
y = -3000-600
y = - 3600
=======================
Cheers,
Stan H.